OpenStudy (abstracthaven):

find the inverse function of f^-1(x)=x+4/3

OpenStudy (ash2326):

y=f(x)=x+4/3 we need f^-1(x) x=y-4/3 so inverse function is f(x)-4/3

myininaya (myininaya):

solve for x and replace x with f(x) and f^{-1}(x) with x

myininaya (myininaya):

\[x=f^{-1}(x)-\frac{4}{3} ...............f(x)=x-\frac{4}{3}\]

OpenStudy (anonymous):

i am going to guess that this could be \[f(x)=\frac{x+4}{3}\]

OpenStudy (anonymous):

in which cases the inverse will be \[f^{-1}(x)=3x-4\]